OpenStudy (anonymous):

Use laplace transforms to solve y’’(t) + 4y(t) = ф(t) with y(0) = 1, y’(0) = 0, where ф(t) = ((4t,0≤t≤1 and 4, t>1))

OpenStudy (anonymous):

you have no idea how we could do this ?

OpenStudy (anonymous):

lol! no! can you give me a start?

OpenStudy (anonymous):

ok the linearity of the laplace transform can help ! L(y"+4y)=L(phi(t))

OpenStudy (anonymous):

still not having the answer ?

OpenStudy (anonymous):

haha! no...

OpenStudy (anonymous):

I will write the answer step by step...when you will be able to continue by yourself...tell me to stop kk ?

OpenStudy (anonymous):

sweet! u will be my lifesaver!!!

OpenStudy (anonymous):

\[L(y"+4y)=L(\phi)\] then \[L(y")+L(4y)=L(\phi)\]

OpenStudy (anonymous):

I guess you can compute L(phi) right .?

OpenStudy (anonymous):

humm...use the formula your just wrote?

OpenStudy (anonymous):

\[L(y")=s^2L(y)-sy'(0)-y(0)\] and \[L(4y)=4L(y)\] so,the equation becomes \[s^2L(y)-1+4L(y)=L(\phi)\]

OpenStudy (anonymous):

still with me .?

OpenStudy (anonymous):

nice...how about this...v tried to solve it...can u check it and see if there are any errors? give me 10 min

OpenStudy (anonymous):

okk that sounds great ! you can send a message to me... if you want to !

OpenStudy (anonymous):

cool!

OpenStudy (anonymous):

I dont see how to attach a file in msg! so...here it is!

OpenStudy (anonymous):

for computing L(phi) you can also use the definition...it's a lot easier :) but you did seems correct ! w8 5 more minutes...to check if there is a miscalculation

OpenStudy (anonymous):

ok take your time!

OpenStudy (anonymous):